# Landauer Principle and Thermodynamics of Computation

### v1

@@ III.3 Landauer Bound in Non-Markovian Process
For the execution of the protocol Lorenzo et al. (2015), the thermalization of the system $\mathcal{S}$ with the environment $\mathbf{E}$ is considered (as shown in Fig. 3). In this model, the environment has N identical noninteracting elements $\mathbf{E}_{N}$, which are conveyed as subenvironments (Fig. 3), and each of them is considered to be in a thermal state $\eta=\otimes_{n=1}^{N}\eta_{n}^{th}$, where $\eta_{n}^{th}=e^{-\beta H_{E_{n}}}/Tr[e^{-\beta H_{E_{n}}}]$. Here, $H_{E_{n}}$ denotes the free Hamiltonian of the n-th subenvironment and $H_{\mathcal{S}}$ denotes the free Hamiltonian of the system. The system will interact with the environment, and the interaction will be a sequence of pairwise collisions with the subenvironments. The environment is of a large size to curtail the situation that the system interacts with the same subenvironment twice. Each collision of this process is described by unitary evolution $U=e^{-igH_{int}\tau}$, where $g$ is the interaction strength, $\tau$ is the collision time, and $H_{int}$ denotes the interaction Hamiltonian of the system and the environment. The information that is stored in the system gets diluted while interacting with the environment. After (n+1) collisions, the state of the system and the environment are respectively described as
@@
$\beta\dot{Q}(t)\geq\dot{S^{\prime}}(\rho),$ (15)
where $S^{\prime}(\rho)=-S(\rho)$. Eq. (15) provides the formulation of the LP in NM dynamics of OQS. The formalism proposed in the work Lorenzo et al. (2015) has twofold power. One, it provides a time-resolved analysis of the erasure by thermalization, which enabled the formulation of the LP for the NM dynamics. Secondly, it was possible to elucidate the role of correlation in the information-erasure processes. This formalism provides the linkage between the erasure process with the intrasystem correlations that arise in the open quantum dynamics for multipartite systems.
The violation of the LB due to strong correlation (proposed in Lorenzo et al. (2015)) is encountered in the spin-1/2 particle system Pezzutto et al. (2016) when used as the framework for the analysis. The sequence of discrete-time collisions of the system with one of the environmental particles at a time occurs in this process, and the Hamiltonian is ruled by the Heisenberg Hamiltonian. In the long time limit, the environment (for non-interacting particles) undergoes homogenization dynamics, i.e., in the long time limit which is equivalent to a large number of collisions of the system with the environment, the state of the system approaches asymptotically to the initial preparation of the environment Ziman et al. (2001); Scarani et al. (2002). If the environment is considered to be composed of interacting particles, the system undergoes an NM process. Here, the elements of the environment are in a thermal state, and in the asymptotic limit homogenization dynamics are encountered. This behavior occurs when the state of the elements of the environment has weak fluctuation. Due to the inter-environment interactions, there is a memory effect in the dynamics of the system, which features the non-Markovianity. It was observed that there is an instantaneous violation of the LB for the system, the cause of which is the strong system-environment correlation.
@@
where $S(\rho||\rho^{\prime})=\text{Tr}\,\rho\,\ln\,\rho-\text{Tr}\,\rho\,\ln\,\rho^{\prime}$ describes the quantum relative entropy, $\rho_{\mathcal{S}_{n}}^{\prime}$ and $\rho_{\mathbf{E_{n}}}^{\prime}$ denotes the marginal states of the system and the environment respectively after the collision. $I(\rho_{\mathcal{S}\mathbf{E_{n}}})$ characterizes the mutual information that quantifies the system-environment correlation based on the intra-collision strength. The second term in Eq. (16) (followed from Eq. (2)) describes the entropy flow from the environment to the system. The LB for the NM process is
@@
where $\Delta\tilde{Q_{n}}=-\Delta Q_{n}$ and $\Delta\tilde{S}_{n}=-\Delta S_{n}$. When $I(\rho_{\mathcal{S}\mathbf{E_{n}}})\leq S(\rho_{\mathcal{S}\mathbf{E_{n}}}^{\prime}||\rho_{\mathcal{S}_{n}}^{\prime}\rho_{\mathbf{E_{n}}}^{th})-S(\rho_{\mathbf{E_{n}}}||\rho_{\mathbf{E_{n}}}^{th}))$, the LP holds as long as the established system-environment correlations are smaller than the upper bound. So, the condition that allows the system to violate the LP is
$I(\rho_{\mathcal{S}\mathbf{E_{n}}})>S(\rho_{\mathcal{S}\mathbf{E_{n}}}^{\prime}||\,\rho_{\mathcal{S}_{n}}^{\prime}\rho_{\mathbf{E_{n}}}^{th})-S(\rho_{\mathbf{E_{n}}}||\,\rho_{\mathbf{E_{n}}}^{th})).$ (18)
It can be generalized, where the system is coupled to a composite environment with $M$ different environments. The LP violation condition for the multiple NM environments is found to be equivalent to that of Eq. (18).
A solution to the violation of the LP has been put forward in the work Zhang et al. (2021). Here, the authors considered a different scenario, the system ($\mathcal{S}$) interacts with $\mathcal{S}_{1}$ (ancillary system) which in turn is coupled to a Markovian environment $\mathbf{E}$. This composite environment induces NM dynamics (Fig. 4). The memory in this model is $\mathcal{S}_{1}$. In the Markovian limit, the conventional LB holds. The modified version of the LB for the heat dissipation that holds for the NM as well as the Markovian regime is
@@
where $\dot{Q}_{\mathcal{S}}(t)$ describes the heat flux from the system to the environment, $\dot{\tilde{S}}(\rho_{\mathcal{S}}(t))$ is the entropy flux from $\mathcal{S}$ to $\mathbf{E}$ ($\tilde{S}(\rho_{\mathcal{S}}(t))=-S(\rho_{\mathcal{S}}(t))$), $\dot{I}(\rho_{\mathcal{S^{\prime}}}(t))$ describes the rate of mutual information that characterizes the correlation between the system and auxiliary system and $\dot{S}(\rho_{\mathcal{S}_{1}}(t)||\rho_{\mathcal{S}_{1}}^{th})$ describes the rate of quantum relative entropy. The generalization of the model for multiple environments has also been reported in this work. The modified LP for the multiple environment case is also found to be valid in both regimes and is a direct consequence of Eq. (2).
Another aspect that was encountered for violation of the LP in the NM domain is the information back-flow. In the work Hu et al. (2022), the authors evaluated the connection between the information back-flow in the NM dynamics and LP. For the analysis, a qubit is considered to be coupled with the environment. If the system is initially in the thermal state, it is inferred from the analysis that there is a one-to-one correspondence between the violation of the LE and the information backflow. Whereas the correspondence does not hold if the initial state of the system has coherence.
@@ IV Landauer limit in Computing
A quantum computer Nielsen and Chuang (2002); Aharonov (1999); DiVincenzo (2000) harnesses quantum mechanical phenomena such as superposition and entanglement Mermin (1990); Linden et al. (2006); Wineland (2013) to perform computations. Qubits, serve as the fundamental units of information in quantum computers and leverage the principles of quantum superposition enabling quantum computers to execute certain types of calculations much more efficiently than classical computers. However, regardless of the computational task, it ultimately needs to be implemented on a physical system, implying that computations are subject to the constraints of the laws of physics. For instance, the LP places a minimum limit on the heat generated during bit erasure, and the quantum speed limit Golub and Ortega (2014); Caneva et al. (2009); Okuyama and Ohzeki (2018); Jones and Kok (2010); Deffner and Campbell (2017); del Campo et al. (2013); Deffner and Lutz (2013) dictates a limitation on how quickly a fundamental logical operation can be carried out. The modern computer uses a binary logic system, and LB for such binary computers is known to be $k_{B}T\ln 2$.
IV.1 Landauer’s Limit For N-based logical computer
A computer using the binary system can be exemplified as a single particle Szilard engine Szilard (1929) where the certainty of the particle’s presence in a chamber corresponds to the recording of 1 bit of information, and the uncertainty in the particle’s location corresponds to the erasure of 1 bit of information. However, computers are not limited to binary logic systems. In principle, they can be based on many-valued logic. For instance, a ternary logic-based computer (“trit”) is founded on the concept of a ternary symmetrical number system and ternary memory element (“flip-flap-flop”) Glusker et al. (2005); Brousentsov (1965); Stakhov (2002); Frieder et al. (1973); Knuth (1998). Recently, computers employing many-valued logic have garnered significance due to fundamental aspects and numerous applications Gottwald and Gottwald (2001); Chang (1958).
The exemplification of the LP for binary logic computers can be supplied by the Brownian particle in a double potential well as shown in Fig. 5. For a symmetrical well and a random bit $a_{0}=a_{1}=\frac{1}{2}$ (where $a_{0}$, $a_{1}$ denotes the probability) the LB is recovered by implementing Eq. (2). Now, whether LL holds for N-based logic is explored in Bormashenko (2019b). In the work, the authors have considered ternary logic, i.e., trit computing element for the analysis which is further generalized for N-based logic. Similar to binary logic computers, the LP is illustrated for ternary logic computers using a Brownian particle in a symmetric triple-well potential. For a random bit, $a_{0}=a_{1}=a_{2}=\frac{1}{3}$ (where $a_{0}$, $a_{1}$, $a_{2}$ denotes the probability), the LB is reached for the ternary logic by implementing Eq. (2), and this result has been further extended to N-based logic.
@@ IV.1 Landauer’s Limit For N-based logical computer
As computational algorithms become more complex, parallel computing becomes essential for the efficient execution of protocols. Parallel computing, as discussed in Kumar et al. (1994); Barney et al. (2010); Melhem (1992); Golub and Ortega (2014), involves the simultaneous utilization of multiple processors to solve algorithms. The primary objective is to distribute the workload among several processors to solve problems more swiftly or handle larger problems within the same timeframe. In the study Konopik et al. (2021), the authors examined the energy cost of finite-time irreversible computing using non-equilibrium thermodynamics principles. It has shown that the energy cost for computational tasks in a parallel computer, within a given finite time, closely adheres to the LL. This cost remains bounded even as the computational problem size increases.
IV.2 Landauer’s bound in the presence of time-symmetric protocol
Investigations into energy dissipation with time-symmetric protocols, demonstrates a fundamental trade-off between computational accuracy and energy cost Riechers et al. (2020). It shows that reducing logical error requires increasing energy dissipation, particularly under nonreciprocal operations, and thus time-symmetric protocols prevent computation from reaching the LB. In a subsequent work Wimsatt et al. (2021), authors utilized thermodynamic analysis of time-symmetric procedures (as prescribed in Riechers et al. (2020)) to thoroughly examine the trade-offs between accuracy and dissipation during information erasure. The authors employed nonequilibrium information thermodynamics to calculate the minimum energy dissipation needed for reliable erasure under time-symmetric control protocols. The energy costs associated with reliable erasure were found to be higher than those implied by the LB on information erasure. Moreover, these costs diverge in the limit of perfect computation. Therefore, the creation of time-asymmetric protocols is deemed necessary for effective and precise thermodynamic computation. Consequently, it is concluded that time asymmetry serves as a crucial design principle for thermodynamically efficient computing, warranting further investigation.
@@ V Experimental Validation of Landauer’s Principle
With our ever-increasing control over the miniaturized scale and the advancement of quantum technology, the validation of the LB is becoming more and more plausible day-by-day. Here in this section, we will focus on the recent experimental advancement to test LP in different technological platforms both in the stochastic and quantum realm.
@@ V.1 Optics Based Technology
The test of LP was possible (that remained untested over five decades) after the two basic advancements. One of them was to develop a method to investigate the work done on the particle as well as the heat dissipation by the particle based on the information on the trajectory of the particle and its potential. It was proposed and tested in the seminal paper by Sekimoto Sekimoto (1997, 2010). The second advancement is the development of methods to impose user-defined potential on small particles. For e.g., the usage of the localized potential force. This potential is created by optical tweezers and is used to explain LP under partial erasure Bérut et al. (2012). In this work, the authors have considered an overdamped colloidal particle inside the double potential well (Fig. 6) which is created by focusing a laser alternatively at two different positions with a high switching rate. The form of the potential well is determined by the intensity of the laser and the distance between the two focal points. If the particle is on the left side of the well, the state of the system is denoted by ‘0’, whereas if the particle is on the right side, it is ‘1’. The experimental process for this method can be summarized as follows:
@@
$\bullet$ Finally, after the bead ends in the required well independent of the initial state (this causes the memory erasure of the process when it is set to 1), the barrier is raised to its previous stage.
@@
To explore the LP for the underdamped and overdamped systems Dago et al. (2021); Dago and Bellon (2022), differential interferometer Paolino et al. (2013) have been considered as the platform for the analysis. The working system is a micromechanical oscillator (the role played by a conductive cantilever in the experiment) confined in a double potential well. Here, the cantilever is the memory cell. Within the framework of stochastic thermodynamics, it has been shown in the work Dago et al. (2021) that one can reach the LB with high precision within a time scale of 100 ms which was previously recorded to be 30 s. The work was further extended by the authors to explore the overdamped and underdamped case within the stochastic thermodynamic framework with fast operation Dago and Bellon (2022). There they encountered a transient temperature rise, so the mean work to erase the information increases but is still bounded by LP.

### v2

@@ III.3 Landauer Bound under Non-Markovian Dynamics
For the execution of the protocol Lorenzo et al. (2015), the thermalization of the system $\mathcal{S}$ with the environment $\mathbf{E}$ is considered (as shown in Fig. 3). In this model, the environment has N identical noninteracting elements $\mathbf{E}_{N}$, which are conveyed as subenvironments (Fig. 3), and each of them is considered to be in a thermal state $\eta=\otimes_{n=1}^{N}\eta_{n}^{th}$, where $\eta_{n}^{th}=e^{-\beta H_{E_{n}}}/Tr[e^{-\beta H_{E_{n}}}]$. Here, $H_{E_{n}}$ denotes the free Hamiltonian of the n-th subenvironment and $H_{\mathcal{S}}$ denotes the free Hamiltonian of the system. The system interact with the environment, and the interaction is governed by a sequence of pairwise collisions with the subenvironments. The environment is of a large size to curtail the situation that the system interacts with the same subenvironment twice. Each collision of this process is described by unitary evolution $U=e^{-igH_{int}\tau}$, where $g$ is the interaction strength, $\tau$ is the interaction time, and $H_{int}$ denotes the time-independent interaction Hamiltonian of the system and the environment. The information that is stored in the system gets diluted while interacting with the environment. After (n+1) collisions, the state of the system and the environment are respectively described as
@@
$\beta\dot{Q}(t)\geq\dot{\tilde{S}}(\rho),$ (15)
where $\tilde{S}(\rho)=-S(\rho)$. Eq. (15) provides the LP in OQS formulation involving heat and entropy fluxes. This relation provides a time-resolved analysis of the erasure by thermalization. Moreover, it is possible to elucidate the role of correlation among the subsystems, contributing through the entropy, in the information-erasure processes. This in turn enable us to probe the effect of the NM dynamics in information-erasure.
The local violation of the LB due to the system-environment correlation (proposed in Lorenzo et al. (2015)) is demonstrated in the spin-1/2 particle system Pezzutto et al. (2016) when interaction between the system and environment is governed by the Heisenberg interaction. The sequence of discrete-time collisions of the system with one of the environmental particles at a time occurs in this process. In the long time limit, the environment (for non-interacting particles) undergoes homogenization dynamics, i.e., in the long time limit which is equivalent to a large number of collisions of the system with the environment, the state of the system approaches asymptotically to the initial preparation of the environment Ziman et al. (2001); Scarani et al. (2002). If the environment is considered to be composed of interacting particles, the system undergoes an NM dynamics. The elements of the environment are in a thermal state, and in the asymptotic limit, homogenization dynamics are encountered. This behavior occurs when the state of the elements of the environment has weak fluctuations. However, due to the inter-environment interactions, there is a memory effect in the dynamics of the system. It has been observed that during the time evolution, the system-environment correlation grows strong enough to exhibit NM features leading to an instantaneous violation of the LB.
@@
where $S(\rho||\rho^{\prime})=\text{Tr}\,[\rho\,\ln\,\rho]-\text{Tr}\,[\rho\,\ln\,\rho^{\prime}]$ describes the quantum relative entropy, $\rho_{\mathcal{S}_{n}}^{\prime}$ and $\rho_{\mathbf{E_{n}}}^{\prime}$ denotes the marginal states of the system and the environment respectively after the collision. $I(\rho_{\mathcal{S}\mathbf{E_{n}}})$ characterizes the mutual information that quantifies the system-environment correlation based on the intra-collision strength. The second term in Eq. (16) describes the entropy flow from the environment to the system. The modified LB for the NM process can be expressed as
@@
where $\Delta\tilde{Q_{n}}=-\Delta Q_{n}$ is the heat dissipated to the environment and $\Delta\tilde{S}_{n}=-\Delta S_{n}$ entropy decrease of the system. When $I(\rho_{\mathcal{S}\mathbf{E_{n}}})\leq S(\rho_{\mathcal{S}\mathbf{E_{n}}}^{\prime}||\rho_{\mathcal{S}_{n}}^{\prime}\rho_{\mathbf{E_{n}}}^{th})-S(\rho_{\mathbf{E_{n}}}||\rho_{\mathbf{E_{n}}}^{th}))$, the LP holds, or we can say that the established system-environment correlations are smaller than the upper bound. Therefore, the condition that allows to violate the LP is
$I(\rho_{\mathcal{S}\mathbf{E_{n}}})>S(\rho_{\mathcal{S}\mathbf{E_{n}}}^{\prime}||\,\rho_{\mathcal{S}_{n}}^{\prime}\rho_{\mathbf{E_{n}}}^{th})-S(\rho_{\mathbf{E_{n}}}||\,\rho_{\mathbf{E_{n}}}^{th}).$ (18)
The authors have demonstrated violation of LP by taking the system-environment interaction to be the Heisenberg interaction. Their findings further suggest that the complexity of the environment has little to do with the violation of LP under NM dynamics. They have further shown that their results can be generalized to where the system is coupled to multiple environments. The LP violation condition for the multiple NM environments is found to be a straightforward generalization of Eq. (18).
A solution to the apparent violation of the LP in the NM domain has been put forward in the work Zhang et al. (2021). Here, the authors considered a different scenario, the system ($\mathcal{S}$) interacts with $\mathcal{S}_{1}$ (ancillary system) which in turn is coupled to a Markovian environment $\mathbf{E}$. This composite environment induces NM dynamics (Fig. 4) through the memory effect induced by $\mathcal{S}_{1}$. In the Markovian limit, the conventional LB holds. The modified version of the LB for the heat dissipation that holds for the NM as well as the Markovian regime is
@@
where $\dot{Q}_{\mathcal{S}}(t)$ describes the heat flux from the system to the environment, $\dot{\tilde{S}}(\rho_{\mathcal{S}}(t))$ is the entropy flux from $\mathcal{S}$ to $\mathbf{E}$ ($\tilde{S}(\rho_{\mathcal{S}}(t))=-S(\rho_{\mathcal{S}}(t))$), $\dot{I}(\rho_{\mathcal{S^{\prime}}}(t))$ describes the rate of mutual information that characterizes the correlation between the system and auxiliary system and $\dot{S}(\rho_{\mathcal{S}_{1}}(t)||\rho_{\mathcal{S}_{1}}^{th})$ describes the rate of quantum relative entropy. Their model allows them to study the Markovian limit, and it has been shown that in the Markovian limit the above equation reduces to (15). The generalization of the model for multiple environments has also been reported in this work. The modified LP for the multiple environment case is also found to be valid in both regimes. Importantly, the quantum information theoretic framework that has been adopted to derive the the modified LBs in the NM domain discussed here broadly belongs to the same framework used to derive Eq. (2).
In Hu et al. (2022), the authors have evaluated the connection between the information back-flow in the NM dynamics and the violation of LP. For the analysis, a qubit is considered to be coupled with the environment. If the system is initially in the thermal state, it is inferred from the analysis that there is a one-to-one correspondence between the violation of the LP and the information backflow. Whereas, the correspondence does not hold if the initial state of the system has coherence.
@@ IV Landauer limit in Computing
A quantum computer Nielsen and Chuang (2002); Aharonov (1999); DiVincenzo (2000) harnesses quantum mechanical resources such as superposition and entanglement Mermin (1990); Linden et al. (2006); Wineland (2013) to perform computations. Qubits serve as the fundamental units of information in quantum computers and leverage the principles of quantum superposition enabling quantum computers to execute certain types of calculations much more efficiently than classical computers. However, regardless of the computational task, it ultimately needs to be implemented on a physical system, implying that computations are subject to the constraints of the laws of physics. For instance, the LP places a minimum limit on the heat generated during a bit erasure, and the quantum speed limit Golub and Ortega (2014); Caneva et al. (2009); Okuyama and Ohzeki (2018); Jones and Kok (2010); Deffner and Campbell (2017); del Campo et al. (2013); Deffner and Lutz (2013) dictates a limitation on how quickly a fundamental logical operation can be carried out.
IV.1 Landauer Limit For N-based logical computer
A computer using the binary logic can be exemplified as a single particle Szilard engine Szilard (1929) where the certainty of the particle’s presence in a chamber among the two corresponds to the recording of 1 bit of information, and the uncertainty in the particle’s location corresponds to the erasure of 1 bit of information. However, computers are not limited to binary logic systems. In principle, they can be based on many-valued logic. For instance, a ternary logic-based computer (“trit”) is founded on the concept of a ternary symmetrical number system and ternary memory element (“flip-flap-flop”) Glusker et al. (2005); Brousentsov (1965); Stakhov (2002); Frieder et al. (1973); Knuth (1998). Recently, computers employing many-valued logic have garnered significance due to both fundamental aspects and numerous applications Gottwald and Gottwald (2001); Chang (1958).
The exemplification of the LP for binary logic computers can be supplied by the Brownian particle in a double potential well as shown in Fig. 5. For a symmetrical well and a completely random bit with $a_{0}=a_{1}=\frac{1}{2}$, where $a_{0}$, $a_{1}$ denotes the probability of being trapped in the wells, the LB is reached. Now, whether LB holds for N-based logic is explored in Bormashenko (2019b), where the authors have considered ternary logic, i.e., trit computing element for the analysis which is further generalized for N-based logic. Similar to binary logic computers, the LP is illustrated for ternary logic computers using a Brownian particle in a symmetric triple-well potential. For a random bit, $a_{0}=a_{1}=a_{2}=\frac{1}{3}$, where $a_{0}$, $a_{1}$, $a_{2}$ denotes the probability of being trapped in the wells, the LB is reached for the ternary logic, and this result has been further extended to N-based logic.
@@ IV.1 Landauer Limit For N-based logical computer
As computational algorithms become more complex, parallel computing becomes essential for the efficient execution of protocols. Parallel computing, as discussed in Kumar et al. (1994); Barney et al. (2010); Melhem (1992); Golub and Ortega (2014), involves the simultaneous utilization of multiple processors to solve algorithms. The primary objective is to distribute the workload among several processors to solve problems more swiftly or handle larger problems within the same time window. In Konopik et al. (2021), the authors examined the energy cost of finite-time computing using non-equilibrium thermodynamics principles. It has been shown that the energy cost for computational tasks in a parallel computer, within a given finite time, closely adheres to the LB. This cost remains bounded even as the computational problem size increases.
IV.2 Landauer Limit in Presence of Time-symmetric Protocol
Investigations into energy dissipation with time-symmetric protocols, demonstrates a fundamental trade-off between computational accuracy and energy cost Riechers et al. (2020). It shows that reducing logical error requires increasing energy dissipation, particularly under nonreciprocal operations, and therefore time-symmetric protocols prevent computation from reaching the LB. In a subsequent work Wimsatt et al. (2021), authors did thermodynamic analysis of time-symmetric procedures (as prescribed in Riechers et al. (2020)) to thoroughly examine the trade-offs between accuracy and dissipation during information erasure. The authors employed nonequilibrium information thermodynamics to calculate the minimum energy dissipation needed for reliable erasure under time-symmetric control protocols. The energy costs associated with this reliable erasure were found to be higher than those implied by the LB. Moreover, these costs diverge in the limit of perfect erasing. Therefore, the creation of time-asymmetric protocols is deemed necessary for effective and precise thermodynamic analysis of computation. Consequently, it is concluded that time asymmetry serves as a crucial design principle for thermodynamically efficient computing, warranting further investigation.
@@ V Experimental Validation of Landauer’s Principle
With our ever-increasing control over the miniaturized scale and the advancement of quantum technology, the validation of the LB is becoming more and more plausible day-by-day. Here in this section, we will focus on the recent experimental advancement to test LP in different technological platforms.
@@ V.1 Optics Based Technology
The test of LP has become possible (that remained untested over five decades) after the two basic advancements. One of them is to develop a method to investigate the work done on the particle as well as the heat dissipation by the particle based on the information on the trajectory of the particle and its potential. It has been proposed and tested in the seminal paper by Sekimoto Sekimoto (1997, 2010). The second advancement is the development of methods to impose user-defined potential on small particles. For e.g., the usage of the localized potential force. This potential is created by optical tweezers and is used to explain LP under partial erasure Bérut et al. (2012). In this work, the authors have considered an overdamped colloidal particle inside the double potential well (Fig. 6) which is created by focusing a laser alternatively at two different positions with a high switching rate. The form of the potential well is determined by the intensity of the laser and the distance between the two focal points. If the particle is on the left side of the well, the state of the system is denoted by ‘0’, whereas if the particle is on the right side, it is ‘1’. The experimental process for this method can be summarized as follows:
@@
$\bullet$ Finally, after the bead ends in the required well independent of the initial state (this causes the memory erasure when it is set to 1), the barrier is raised to its previous stage.
@@
To explore the LP for the underdamped and overdamped systems Dago et al. (2021); Dago and Bellon (2022), differential interferometer Paolino et al. (2013) has been considered as the platform for the analysis. The working system is a micromechanical oscillator (the role played by a conductive cantilever in the experiment) confined in a double potential well. Here, the cantilever is the memory cell. Within the framework of stochastic thermodynamics, it has been shown in the work Dago et al. (2021) that one can reach the LB with high precision within a time scale of 100 ms, which was previously recorded to be 30 s. The work was further extended by the authors to explore the overdamped and underdamped case within the stochastic thermodynamic framework with fast operation Dago and Bellon (2022). There they encountered a transient temperature rise, so the mean work to erase the information increases but is still bounded by LP.